OpenStudy (anonymous):

Which of the following points lie in the solution set to the following system of inequalities? y > -3x + 3 y > x + 2 A. (2,-5) B. (-2,5) C. (2,5) D. (-2,-5)

OpenStudy (ajprincess):

Let us check with the points (2,-5) if they satisfy the inequalities. -3(2)+3=-6+3=-3 -3>-5 so this is nt a solution. let us check with the points (-2,5) -3(-2)+3=6+3=9 9>5 this too is nt a solution. let us chck with (2,5) -3(2)+3=-6+3=-3 5>-3 First inequality is satisfied. Let us see if it satisfies the second inequality too. 2+2=4 5>4 therefore (2,5) is the solution.